OpenStudy (anonymous):

solve for x. a^2x^2+2ax+1=0 but "a" CANT =0

OpenStudy (anonymous):

Make "a" a number that simplifies the problem

OpenStudy (anonymous):

Try a=1

OpenStudy (mertsj):

Use the quadratic formula. a=a^2 b=2a c=1

OpenStudy (anonymous):

a^2x^2 = (ax)^2, so (ax)^2 + 2ax + 1 could be factored as (ax + 1) (ax + 1)

OpenStudy (mertsj):

\[x=\frac{-2a \pm \sqrt{(-2a)^2-4(a^2)(1)}}{2(a^2)}\]

OpenStudy (mertsj):

\[x=\frac{-2a \pm \sqrt{0}}{2a^2}\]

OpenStudy (mertsj):

\[x=\frac{-2a}{2a^2}=-\frac{1}{a}\]